13x+7y=88
x+8y=-38
Solve this problem using the Solution method

Answer by mananth(16949)

13x+7y = 88 .............1
x+8y=-38 .............2
Eliminate y
multiply (1)by -8
Multiply (2) by 7
-104x-56y=-704
7x+56y =-266
Add the two equations
-97x=-970
/-97
x=10
plug value of x in (1)
13 x + 7 y = 88
130 + 7 y = 88
7 y = 88 -130
7 y = -42
y = -6
(10,-6)